OpenStudy (anonymous):

Solve each equations for x 9-(6x-8)=-8(6x-8)

OpenStudy (anonymous):

set (6x-8) = t (just a random variable, could be anything (this is not neccesary but it makes it easier) so you get 9 - t = -8t if you solve this you get 9 = -7t t = 9/-7 then you fill the 6x-8 back in. so you get 6x-8 = 9/-7 now solve for x